OpenStudy (anonymous):

x^2+y^2-2x+6y+9=0 . What is the center of the circle of this equation? (:

OpenStudy (callisto):

x^2+y^2+Dx+Ey+F=0 Centre = ( -D/2 , -E/2)

OpenStudy (mimi_x3):

\[x^2-2x+y^2+6y+9=0\] Compete the square. \[(x-1)^2-(1)^2+(y+3)^2-(3)^2+9=0\] \[(x-1)^2+(y+3)^2-1=0\] \[(x-1)^2+(y+3)^2=1\] Centre \(a,b\) and radius \(r\) units is: \[(x-a)^2+(y-b)^2=r^2\] soo..centre = 1 and -3

OpenStudy (anonymous):

thank you everyone! :D